OpenStudy (anonymous):

Use the substitution method to solve the system of equations. x + 2y = 15 2x – 3y = 2

OpenStudy (anonymous):

x=15-2y

OpenStudy (anonymous):

Answers are A (7, 4) B (2, 4) C (4, 7) D (4, 2)

OpenStudy (anonymous):

7,4

OpenStudy (lalaly):

from first equation x=15-2y substitute in second equation 2(15-2y)-3y=2 30-4y-3y=2 28=7y y=4 x=15-2(4)= 15-8= 7 so (7,4)